OpenStudy (anonymous):

For the pair of functions, find the indicated sum, difference, product, or quotient. f(x) = square root 3x + 5, g(x) = square root25x - 16 Find (fg)(x).

OpenStudy (anonymous):

\[f[x]=\sqrt{3 x+1} , g[x]=\sqrt{25x-16} \text{ ?} \]

OpenStudy (anonymous):

Sorry,\[f[x]=\sqrt{3 x+5}\text{ }? \]

OpenStudy (anonymous):

A Mathematica solution with comments is attached.

OpenStudy (anonymous):

u cant just multiply f(x) & g(x) coz if 3x+1 & 25x-16 are negative...there r two conditions-1)f & g r both positive multiply directly ,that wud be first answer and 2) in the second condition put sq(3x+1)=X1+Yi & sq(25x-16)=X2+Y2i,insert this value and find the second solution

OpenStudy (anonymous):

I took the radical signs as positive as you see. I don't see where any equations are involved here. could you go ahead with your own solution. Thank you.

OpenStudy (anonymous):

thank u